OpenStudy (anonymous):

how to write a formula for the statement the probability that John owns a dog given that Sally own a dog is 0.6...

OpenStudy (kropot72):

Let J be the event 'John owns a dog' and let S be the event 'Sally owns a dog'. The probability that John owns a dog given that Sally owns a dog can be written as: \[\large P(J|S)\]

OpenStudy (anonymous):

And then you would P(J/S)=P(A intercept B)/S correct?

OpenStudy (kropot72):

\[\large P(J|S)=\frac{P(J \cap S)}{P(S)}=0.6\]

OpenStudy (anonymous):

Thank you.

OpenStudy (kropot72):

You're welcome :)